UT 20121029 AO#1

Observer: D.Miller (LBTO),K.Boutsia(INAF) JHill remote(Tucson), O.Kuhn remote (Tucson)
Telescope Operator: Geno.B.(LBTO)
SW Support: MDLPena (Tucson)
AO Support: DMiller (LBTO)
Telescope Support: Kevin N (LBTO)
Instrument Support: Elliot S (LBTO)

Summary

We successfully sent a fully automatic binocular adaptive ace preset, along with a runao and the result was closed loop AO on both sides of the telescope.

This required typing 5 IDL commands and from preset to closed loop AO took about 12 minutes. There was no manual interaction, just the requests the science instruments will use.

Note: we spent about 30 minutes previous to the request for binocular closed loop manually adjusting the pointing model, collimating to remove some aberrations, etc, so the 12 minutes has be taken carefully. The 30 minutes spent adjusting because of problems can/should/must be solve. Some of the problems are general telescope problems such as pointing and collimation and primary jumps. However, there are couple of problems with the binocular preset logic.

Needless to say, we (I) was excited to have the both sides in Closed Loop in an "automatic" procedure.

Other items:

  • John and Tina confirmed that LBC now guides properly with 2012H
  • Alfio fixed several problems on the SX AO and we closed the loop with 2012H
  • Dave collected LUCI data
  • Binocular operation
    • Able to preset in binocular mode (take some manual intervention)
    • Found several problems with binocular setup
    • Tested and confirmed the proper operation of Olga's Binocular Range Balance
    • Closed the AO loops on both side simultaneously
  • Tested the upgraded Center-of-weight GCS (phase 1 on the way to Source Extractor)
    • Collimate properly straight away.
    • The upgraded COW seems more stable than old COW
    • Found a problem with off-axis calculation in GCS
    • Ran with new GCS most of the night to gather robustness information
    • Will put old GCS back in place for science operation
    • Found problem with old wfsc image chop routine in GCS that limit Y value of pupil hotspot to < 290
  • Improve Active Optics parameters at LFBG and RFBG
    • Determined parameters yesterday
    • Installed and tested tonight
    • Will leave in new parameters, with old GCS, for science operation
    • Tested old GCS with new parameters and parameter file. OK
  • Verified DX AO operation with 2012H: Complete
    • Camera lens could not converge problem about 1/4th of the preset failed.
    • Several minor problems found, most fixed by Alfio on the fly
    • Closed loop at all bin levels and TTM work
    • AO offsets were successful
    • Could not communicate with the IRTC via the "acq" routines that collect Loop and IRTC data. Said "IRS not running"
  • Verified SX AO operation with 2012H: Not Complete
    • Camera lens could not converge problem about 6/7th of the preset failed.
      • time spent trying to move lens should be smaller. Now is several minutes. Should fail more quickly so can resend with different rotator angle
    • Big focus difference between AO pyramid and the AGw Shack-Hartmann
    • Only closed the AO loop 2 out of 11 tries.
      • 7 failures were "camera lens failed to converge"
      • 2 failures were 'ERROR_VALUE_OUT_OF_RANGE' due to (we think) high background from moon and impending dawn
      • 2 successes were in binocular operation (at the same rotator angle)
    • Did not test AO offsets

==> Discussed with Alfio the possibility that the dichroic in front of LUCI is at a different angle than the IRTC and this is causing the pupil wobble on the SX side. Similarly, could a difference between the PISCES dichroic and the IRTC dichroic cause the DX pupil wobble?

==> Note: It worked very well to have Alfio and Marco on Skype during our nights of AO operation. When a problem or error was encountered they usually could find and fix the problem in a few minutes and we could immediately confirm the fix on-sky.

Details

Start the night with LBC guiding tests*

01:30 Slew to NGC6819 to test LBC blue
  • dofpia,/X2,/Blueonly

01:38 Start the OB with the dither pattern - N6819_test

02:03 Done with LBC tests: John claims victory

Binocular Tests

02:27 Reconfig done, Geno

03:20 preset AO869, MODE=ACTIVE, GS=7, PA=-90 gs_mag = 12.65 (no vignetting)

03:20 First new GCS image = left_wfscimage00056.fits Very nearly the same zernikes. I had centering on. Turned it off in IDL and zernikes are exactly the same, with +-1 nm on a couple zernikes.

03:30 preset AO869, MODE=ACTIVE, GS=7, PA=-90 gs_mag = 15.79 (no vignetting)

04:00 preset AO869, MODE=ACTIVE, GS=7, PA=-90 gs_mag = 15.98 (slight vignetting)

04:04 getHotspot = 225, 210 x,y -450 -1000 04:05 setHotspot = 220, 210 image 115 x,y -200 -700 04:07 setHotspot = 215, 210 image 118 x,y 100 -600 04:08 etHotspot = 220, 200 image 11 x,y lost guiding

04:09 left_acquire000013.fits acquistion failed but could see star. There is a gradient.

04:15 Authorize LUCIFER and IRTC for binocular tests

04:20 Troubles with IDL binocular preset routines, plus no guide stars found. Go to monocular mode on each side and Geno set IE, CA with spiral search.

05:00 pointing set with M5 star

04:55 Had to change file /lbt/tcs/2012H/aos/etc/aos.conf AOSLrotatorIndex from 2 (BBG) to 0 (FBG). Now pupil rerotator follows instrument rotator properly.

05:05 Alfio fix a logic problem when an instrument is authorized and the WUnit is in RefAcquired mode. Restarted AOSL and tested. Worked.

05:09 binocular preset in acuire mode on M5 star. Star found on left, but not right. Geno found that the IE, CA on right were not what he had determined and set earlier. He set them to the proper value and the star appeared in guide acquire image

05:10 binocular preset in active mode on BS9188. Stars seen in acquire image on both sides, but left had a lot of Z6 and Z4 so guiding failed. Send monocular acquire preset to left and adjust collimation by hand. During this time, right was guiding and collimating.

05:30 Send monocular preset to left, guide and collimate.

05:40 Fully collimated binocular telescope performed with fully binocular IDL preset.

LUCI data acquisition

05:40 Turn the telescope over to Dave to collect LUCI information

Binocular Tests

06:18 Authorize IRTC and LUCI

06:22 binoc Preset to BS9188. Star found, guide and act opt on LEFT. No star found on right. IE and CA had changed. Geno set them back to correct values.

06:24 monoc preset on RIGHT. Had accidentally sent a

06:32 Fully collimated on both sides. FWHW ~0.6 on both. Elevation = 60.

06:33 Olga send Binocular Range Balance request. Success. Change DX by about 10" (at least IE/CA). Had some troubles guiding again on Left. Looked like star was still moving when guide image was taken (long tail).

06:40 Send another BRB. Success. Change about 3" in IE. Probable due to small change in elevation.

06:48 Send another BRB. Success. Change about 1.5" in IE.

06:53 Send binoc preset. Failed. Problem that each guider request a change in mount position to put the star on the hotspot. This results in the stars out from the center of the

left_acquire000021.fits right_acquire000006.fits

left_guiderimage001408.fits right_guiderimage000231.fits

07:07 Send BRB. Change IE by 7"

07:15 BinPreset to BS9107 El 75. Failed. No star in right acquisition. Left saw it but failed.

07:20 Try spiral search on right. Star not found.

07:32 BinPreset to M5_1133 in ACQUIRE mode. Star on LEFT, not on RIGHT. Back to IE/CA on RIGHT that had for BS9188. Star just off the left side. Geno center.

07:36 BinPreset ACTIVE to BS9107. Failed. Stars were off from hotspot. Mount movement problem during acquisition.

07:30 BinPreset ACQUIRE to BS9107

07:47 (about) we were collimating in binocular mode then the we lost the star on LEFT from guider. Olga saw the LEFT primary mirror X total collimation value of -2.9. When we preset back to the same star the primary mirror X total collimation value wast -0.4

08:00 BRB move about -1.59"

08:03 RBR moves of order 0.1"

08:20 Bin runao.

08:27 Binocular Closed Loop AO!!

08:34 Alfio found that the flux on the left down by 25% (1500 photon/pixel) compared to the right (2000 photon/pixel). Caused different binning (3/2) and speed (199/400). Alfio is looking into it.

08:42 StopAO both side successful. guiding and act opt start

Verify DX AO with 2012H

08:50 Authorize IRTC on RIGHT only

08:56 runao failed with "camera controller in the wrong state". Could not get WFS Control gui to respond. Restarted wfsarb. Tried to power off, but wfsarb said already in power off state. Turned power off at HW gui. w_stop, w_start and power on system.

09:15 preset AO198, ace. Mag = 8.5, bin1. Success

09:24 offset 2, 0 success

09:24 offset, 0, 2 success

09:24 offset, 90, 0 success loop paused

09:26 offset, 0,0 success loop resumed

09:36 cannot save IRTC and loop date with acq routine on wfsdx. Say IRS is not running

09:39 preset AO1078, ACE, Mag = 11.8, No star seen on CCD47. Bad coords

09:46 preset AO956, ACE, Mag = 11.7, Star found. Guide star double but OK

09:49 runao Failed camera lens could not converge

09:53 preset AO956, ACE with different guide star and PA. runao failed because mag too bright. Catalog said 11.7 R, measured 7.2 on CCD39

10:02 preset AO958, mag = 11.1 seeing = 0.9 FWHM EL = 50 Success

10:14 preset AO337, mag = 13.2, EL = 30 failed no guide star found

10:15 preset to M5 star, Geno align

10:15 preset AO337, mag = 13.2, EL = 30 Failed, camera lens could not converge

10:23 preset AO337, different guide star, PA=180, bin2 success

10:31 preset AO236, mag = 15.2, EL = 60 Failed camera lens failed to converge

10:35 Roll back to old GCS

10:45 preset AO236, TTM, different GS and PA: IDL FIle error at 10:49:52 Could not find in SetRecMat. Alfio looking in to it. Found error in LUT. Try TTM again in a bit

10:50 preset AO236, ACE Success

11:00 preset AO236, TTM success

Verify SX AO with 2012H

11:15 Authorize LUCIFER on SX

11:17 preset AO236, TTM success: Failed camera lens did not converge

11:24 preset AO236 with different PA. Success

11:30 preset AO969 ACE, EL=30, mag = 11.7 - fwhm on GCS~0.9". Failed camera lens failed to converge

11:44 preset AO969 ACE, different GS and PA: Success

11:52 preset AO307 ACE, EL=68, mag = 8.4, star found but completely out of focus on the GCS

11:58 guider is loosing the star - the cross is someplace and the star in another - this causes big jumps in guiding - ROLL BACK GCS

12:02 repeat the same preset with new GCS - preset AO307

12:05 the same thing happens with the guider.... this time we actually lost the star!!!!!

12:06 preset AO307 with different GS and PA Failed camera lens failed to converge

12:13 preset AO307 with different GS and PA Failed - AcquireRef[88] failed - 'ERROR_VALUE_OUT_OF_RANGE'
  • strange because it is the same star as before and it only gave the camera lens error

12:20 repeat preset AO307 - same error

12:23 preset AO308, ACE, EL=67 - same problem with jumping guider - lost the star

12:27 preset AO308 with different GS and PA - Failed AcquireRef[96] failed - 'ERROR_VALUE_OUT_OF_RANGE'
  • probably the backgournd is too high and the AO system gets confused

12:33 preset AO326, ACE, EL=64, failed - camera lens failed to converge

12:40 preset AO326, ACE, different GS and PA - failed - camera lens failed to converge

12:47 preset AO326, ACE, same GS different PA - failed - camera lens failed to converge

12:55 Close dome

==> Discussed with Alfio the possibility that the angle of the dichroic in front of LUCI is at a different angle than the IRTC and this is causing the pupil wobble on the SX side. Similarly, could a difference between the PISCES dichroic and the IRTC dichroic cause the DX pupil wobble?

-- KonstantinaBoutsia - 29 Oct 2012

Topic revision: r8 - 30 Oct 2012, LeeDettmann
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback